# Unveiling Hallucination in Text, Image, Video, and Audio Foundation Models: A Comprehensive Review

### v1

@@ 8 Future Directions
Researchers are actively investigating hallucination mitigation techniques as the challenge of hallucination can be crucial in sensitive areas where generating fictional or incorrect content could have serious consequences Tonmoy et al. (2024)Rawte et al. (2023b). Here are the potential directions for addressing this critical issue of hallucination in these FMs:
Data Resources: Recent studies have highlighted the efficacy of simple fine-tuning on carefully curated high-quality samples for reducing hallucinations, surpassing the impact of large-scale fine-tuning and reinforcement learning approaches. For knowledge-intensive domains, the development of entity-centered fine-tuned instructions that integrate structured knowledge from knowledge graphs shows promise in enhancing accuracy and relevance. Additionally, employing alignment techniques tailored to specific tasks or fields has proven effective in mitigating hallucinations. As research in this area progresses, more resources focused on improving alignment through task-specific or domain-adapted approaches are expected, further bolstering the reliability of language models in generating factual and trustworthy content.
Automated Evaluation: Developing specialized evaluation metrics that consider factors such as factual accuracy and coherence can be useful for hallucination detection. Combining automated evaluation with human judgments via crowdsourcing can capture nuanced aspects challenging for automated systems alone. Additionally, adversarial testing methodologies are also being developed to expose AI systems to crafted inputs, aiding in identifying weaknesses and enhancing resilience against hallucination. Furthermore, fine-tuning FMs on datasets emphasizing fact-checking and accuracy offers another avenue to improve content reliability and reduce the occurrence of hallucinations.
Improving Detection and Mitigation techniques: Mitigating hallucinations in FMs necessitates a multifaceted approach that leverages reasoning mechanisms, knowledge graph integration, specialized fact-checking models, bias mitigation techniques, and active learning methodologies. Emerging techniques like Chain of Thought (CoT) Wei et al. (2022) and Tree of Thought (ToT) Yao et al. (2024) bolster these models’ reasoning capabilities, potentially reducing hallucinations. Integrating knowledge graphs enhances understanding of factual information and concept relationships, aiding content generation and fact-checking. Specialized verification models cross-reference outputs against curated knowledge to identify inaccuracies, while bias detection and mitigation techniques promote fairness. Finally, ethical guidelines and regulatory frameworks governing the responsible use of curated knowledge in AI development mitigate risks and foster public trust, collectively improving the quality, accuracy, and trustworthiness of AI-generated content.
Multimodal Hallucination: Addressing hallucinations in multimodal large Foundation models requires a comprehensive approach spanning data-centric initiatives, cross-modal alignment efforts, architectural innovations, standardized benchmarking, reframing of hallucination, and enhancing interpretability and trust. Data-centric techniques for robust data collection, augmentation, and calibration ensure diverse and high-quality training data. Cross-modal alignment focuses on aligning representations across modalities through sophisticated architectures. Model architectural advancements involve designing specialized models capable of handling complex linguistic and visual inputs effectively. Establishing unified metrics and standardized benchmarks enables accurate assessment of hallucination and reliable performance evaluations. Reframing hallucination as a feature explores its integration into downstream applications, optimizing for the human experience. Finally, developing techniques for interpreting model behavior, visualizing internals, and improving reliability assessment fosters trust in MLLMs. This multifaceted approach collectively addresses the critical hallucination challenge, paving the way for more reliable and trustworthy multimodal AI systems.
@@ 9 Conclusion
This survey paper systematically categorizes existing research on hallucination within FMs, providing comprehensive insights into critical facets, including detection, mitigation, tasks, datasets, and evaluation metrics. It addresses the pervasive impact of hallucination in FMs, acknowledging its impact across diverse domains. By examining recent advancements in detection and mitigation techniques, the paper underscores the importance of addressing this challenge, given FMs’ indispensable role in critical tasks. Its primary contribution lies in introducing a structured taxonomy for classifying hallucination in FMs, spanning text, image, video, and audio domains.

### v4

@@ 6 Future Directions
Researchers are actively investigating techniques to mitigate hallucinations, which is crucial for sensitive applications Tonmoy et al. (2024)Rawte et al. (2023). The Key future directions can include:
Data Resources: Fine-tuning carefully curated high-quality data, integrating structured knowledge from knowledge graphs, and employing task/domain-specific alignment techniques to enhance accuracy and relevance.
Automated Evaluation: Developing specialized metrics for factual accuracy and coherence, combining automated evaluation with human judgments, adversarial testing to identify weaknesses, and fine-tuning fact-checking datasets.
Improving Detection and Mitigation: Leveraging reasoning mechanisms (e.g., Chain of Thought Wei et al. (2022), Tree of Thought Yao et al. (2024)), knowledge graph integration, specialized fact-checking models, bias mitigation techniques, active learning methodologies, and ethical guidelines/regulatory frameworks.
Multimodal Hallucination: Data-centric initiatives, cross-modal alignment, architectural innovations, standardized benchmarking, reframing hallucination as a feature, and enhancing interpretability and trust for reliable multimodal AI systems.
@@ 7 Conclusion
This survey paper systematically categorizes existing research on hallucination within FMs, providing comprehensive insights into critical aspects such as detection, mitigation, tasks, datasets, and evaluation metrics. It addresses the pressing issue of hallucination in FMs, acknowledging its widespread impact across various domains. The paper underscores the importance of addressing this challenge by examining recent advancements in detection and mitigation techniques, given FMs’ indispensable role in critical tasks. Its primary contribution is presenting a structured taxonomy for classifying hallucination in FMs, spanning text, image, video, and audio domains.
8 Limitation
Previous survey papers primarily focused on hallucination in large language models and did not extensively cover hallucinations in vision, audio, and video modalities. This survey paper aims to provide a comprehensive overview of hallucinations across all modalities, considering that hallucinations can occur in any large foundation model. Despite our efforts to provide a comprehensive summary of recent advancements related to hallucination techniques in all foundational models, we acknowledge that we may miss some relevant work in the field and have covered the papers till May 2024.